OpenStudy (anonymous):

4x²+7x+1 solve using the quadratic formula

OpenStudy (luigi0210):

\[\frac{ -b+/-\sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

is the formula

OpenStudy (rulnick):

a=4, b=7, c=1 in the formula will produce two value (one where the "+/-" is "+" and one where it is "-")

OpenStudy (rulnick):

... and note that you are presumably solving 4x^2+7x+1 = 0.